# Representative electricity price profiles for European day-ahead and intraday spot markets

### v1

@@ Abstract
We propose a method to construct representative price profiles of the day-ahead (DA) and the intraday (ID) electricity spot markets and use this method to provide examples of ready-to-use price data sets. In contrast to common scenario generation approaches, the method is deterministic and relies on a small number of degrees of freedom, with the aim to be well-defined and easy to use. We thereby target an enhanced comparability of future research studies on demand-side management and energy cost optimization. We construct the price profiles based on historical time series from the spot markets of interest, e.g., European Power Exchange (EPEX) spot. To this end, we extract key price components from the data while also accounting for known dominant mechanisms in the price variation. Further, the method is able to preserve key statistical features of the historical data (e.g., mean and standard deviation) when constructing the benchmark profile. Finally, our approach ensures comparability of ID and DA price profiles by design, as their cumulative (integral) price can be made identical if needed.
@@ 1   Introduction
Optimization of energy costs has long been acknowledged as a topic of interest for the economic design and operation of chemical processes and energy systems (Friedler,, 2010). Demand-side management (DSM) and demand response (DR) are energy-flexible operation paradigms that have have raised interest of industrial consumers to actively participate in the energy spot markets by adjustment of electricity demand (Pinson et al.,, 2014; Mitsos et al.,, 2018). With electricity prices on European DA spot markets having reached temporary peak values as high as 4 000 €/MWh (Nord Pool,, 2023), the importance of industrial DSM is expected to grow further in the near future. Thus, economic studies on planning and scheduling of industrial processes necessitate a systematic way to select electricity price scenarios. On the other hand, the process of selecting a price scenario should not distract from the main purpose of the paper. Herein, we thus propose a simple and intuitive approach to determine representative price profiles.
A common approach to select an electricity price profile is the use of a recent historical period, e.g., 24 hours, that exhibits a typical behavior in terms of the price range and transient pattern (e.g., Dalle Ave et al., (2019)). Especially when design and planning studies are performed, multiple historical price profiles can be considered to capture the effect of different fluctuation patterns (e.g., Fürsch et al., (2013)). Long horizons of historical data (e.g., up to one year) may be applied if needed (Meese et al.,, 2016). However, most commonly single-day or single-week price profiles are utilized. The selection of historical prices usually follows a non-systematic procedure, wherefore the representativity and generalizability of case study results from the literature is frequently obscure.
One way to improve the representative character of scenarios is to construct average price profiles over a historical period that respects fundamental characteristics of the spot market price profiles. Knittel & Roberts, (2005) identify electricity price trends on different time scales highlighting recurring patterns and characteristic values of the price time series. Rahimiyan & Baringo, (2016) perform a statistical analysis on historical DA and ID price data to quantify the price correlations. Schäfer et al., (2020) and Germscheid et al., (2022) utilize Fourier transforms to analyze principal frequencies of historical DA prices and ID-DA price deviations, respectively, identifying dominant modes at periods of 0.5 day and 1 day. By averaging one year of historical DA and ID price data, significant time-series frequencies, such as daily or weekly, are taken into account to construct a single profile. These periods correspond to significant price patterns and price fluctuations that can be exploited by most processes. This method, though straightforward, may over-smooth the data and potentially obscure certain less prominent trends.
A collection of more sophisticated approaches to determine representative periods in time-series data is presented in the review by Hoffmann et al., (2020). Clustering historical data provides a set of profiles that represents the entity of the historical data, e.g., by selecting the centroid of each cluster (Poncelet et al.,, 2017). Teichgraeber & Brandt, (2018, 2019) employ clustering to generate representative single-day profiles of one-year DA price data. Although the method reduces the number of scenarios, compared to considering the full collection of available data, multiple optimization problems need to be solved and the results are only optimal with respect to the chosen clustering criteria.
Finally, we mention probabilistic forecasting techniques, which exploit correlations of predicted parameters (e.g., weather forecasts) and are widely used to explain price behavior (Nowotarski & Weron,, 2018). Following Monteiro et al., (2018) and Weron, (2014), these methods can be categorized into prediction interval, density, and threshold forecasting. Univariate and multivariate probabilistic models (Cuaresma et al.,, 2004; Rahimiyan & Baringo,, 2016) are prevalent electricity price forecasting methods, but deep learning approaches like normalizing flows (e.g., Cramer et al., (2023)) are also gaining attention. Electricity prices forecasts enable DR computations in future applications, offering advantages over historical data, especially for long horizons and volatile markets (Nowotarski & Weron,, 2018). Unlike data-averaging techniques, price forecasting methods are not restricted to representing data of the present and recent past. Instead, they can incorporate trends of the future, although commonly restricted to the use of historical data and small forecasting windows (e.g., Alonso et al., (2016); Shah et al., (2021)). Notably, however, both historical data based as well as complex scenario forecasting methods have been rather inaccurate at predicting the long-term evolution of electricy prices (Zareipour et al.,, 2010; Weron,, 2014; Gabrielli et al.,, 2022), since this evolution is subject to policy changes and political events (Yang et al.,, 2017).
@@
Publications on DSM of industrial processes, including our own studies, such as Caspari et al., (2019); Mucci et al., (2023); El Wajeh et al., (2024), heavily rely on the chosen electricity price profiles that must represent current trends and allow for generalizable conclusions. Selecting a single historical profile or averaging historical data favors simplicity. Clearly, averaged price trends do not fully capture all features of historical profiles, as price averaging yields a single and exclusive price scenario. Additionally, data averaging does not account for the temporal evolution of prices to consider future price trends. Nonetheless, simple and sufficiently representative price profiles are usually adequate for evaluation of the theoretical potential of DR applications. Conversely, sophisticated scenario generation methods provide multiple price scenarios, valuable for active market participation. Yet, these methods involve a more complex implementation, while forecasts of the long-term price trends are still subject to high uncertainty. Moreover, an extensive evaluation of all possible price scenarios, product demand situations, and other possible operating disturbances quickly results in a curse of dimensionality and is thus often beyond the scope of research studies on DR.
@@
Herein, we aim to establish a simple method to construct meaningful and representative DA and ID benchmark price profiles to be used within future economic studies, e.g., cost optimal planning and scheduling, power-to-X, and energy-flexible design and operation. Within this method, we desire an option to ensure comparability between the DA and ID prices, which is useful for DR across multiple electricity markets, e.g., in bidding strategies (Silva et al.,, 2022), or scheduling and control (Caspari et al.,, 2020). This property is not provided by existing methods, e.g., Maciejowska et al., (2019); Teichgraeber & Brandt, (2019). The developed price profile will exhibit the following essential properties:
@@
    Comparability of DA and ID price profiles,
@@
To achieve the desired properties, we build our method on previous works on constructing average price profiles (Schäfer et al.,, 2020; Germscheid et al.,, 2022) and introduce adaptations tailored to the applications investigated. Because the focus of DR studies is on the economic potential of electricity price fluctuations rather than steady-state optimization, the price variance over the investigation period is assumed to hold greatest importance for optimal economic decision-making. Thus, our method allows for adjusting the magnitude of fluctuations in the price profile. Additionally, we ensure comparability between the DA and ID scenarios based on the near-zero mean distribution of the ID-DA deviation by equalizing the integrated DA and ID prices over the designated period. To underpin this approach, we provide statistical analyses of hourly DA and quarter-hourly real-time ID prices in historical data from European spot markets. Our approach can extend to other markets and price granularities (e.g., half-hourly) where similar observations hold.
@@
In the remainder of this article, we first introduce a method for constructing a representative 24-hour DA price profile from historical data of the last available 365 days (Section 2.1). Next, we utilize the same historical data to extract the ID price profile, based on ID - DA price deviations (Section 2.2). Then, we introduce a few modifications and apply the same approach to generate single-week (Section 3) representative profiles. Last, we create ready-to-use benchmark DA and ID price profiles for future studies on DSM on the different time-scales presented (Section 4). This open-access data set is also provided via Git (Link).
@@ 2.1 DA price profile
We begin with the construction of the benchmark DA profile of one representative day. To this end, we select a price data set of 365 consecutive days to exclude seasonal effects (Knittel & Roberts,, 2005). The 365 days horizon may be chosen arbitrarily and if desired extended to multiple years. We remark that incidents and trends contained in the price data set, e.g., caused by political events and developments, will affect the price scenario created. Here, we use up-to-date price values of the full year 2023. Similar to our prior works (Schäfer et al.,, 2020; Germscheid et al.,, 2022), we first average the historical DA prices for every day hour, $k=1,2,...,24$, individually111 Notice that for $k\geq 2$, Eq. (1) may be equivalently rewritted as $\overline{DA}_{k}=\overline{DA}_{k-1}+\overline{\Delta DA}_{k}$, where $\overline{\Delta DA}_{k}=1/N\sum_{i=1}^{N}{(DA_{i,k}-DA_{i,k-1})}$. Hence, this profile captures the moving average price development over a day. :
@@
where $N=365$ is the number of days of selected historical data, $DA_{i,k}$ is the DA price at hour $k$ of day $i=1,2,...,N$, and $\overline{DA}_{k}$ is the average DA price of hour $k$. If desired, days with significant outliers, weekends or holidays may be excluded. However, here we consider the full year.
@@
Fig. 1: Average and scaled average ($\beta=1.47$) 24-hour DA price profile generated from historical 2023 data (EPEX,, 2023).
@@
Clearly, setting $\beta=1$ corresponds to no scaling. The resulting scenario exhibits low deviating behavior according to the historical data (Fig. 2). To minimize scenario complexity, we can thus use the $\beta=1$-profile as a moderate scenario. Importantly, the mean value $\overline{DA}$ remains identical irrespective of scaling.
As the daily standard deviation of the average profile might not represent the daily standard deviation of all data due to smoothing of the data (see Schäfer et al., (2020)), we use $\beta$ to correct the averaged results. Thus, in order to obtain a nominal benchmark profile, we choose $\beta$ such that:
@@
For example, for the full year 2023 and EPEX spot market, we receive the value $\beta=1.47$. The resulting DA profile is illustrated by Fig. 1. Here, the characteristic double peak profile is visible with minimal values occurring during the 4th and 14th hour as well as maximal values during the 8th and the 9th. The absolute price range after scaling ($\beta=1.47$) is $103.46\,\mathrm{€/MWh}$ with a mean value of $\overline{DA}=95.18\,\mathrm{€/MWh}$.
If a more distinct price variation is desired, we can specify a larger value of $\beta$. In order to systematically choose a larger $\beta$ factor that corresponds to some extreme behavior of the historical data (Requirement 2), we represent the density distribution of the standard deviations $\sigma(DA_{i,1},...,DA_{i,24})$ across all days $i=1,2,...,N$ through a histogram, Fig. 2. Subsequently, we identify a quantile indicative of high fluctuations within the historical data. We exemplary show the quantile 3.4 ($Q_{3.4}$), corresponding to the 15% highest values of the daily standard deviations, Fig. 2. The value of the quantile is the corresponding standard deviation, namely:
@@
giving $\beta=1.97$ for the full year 2023. Other options could include the choice of a preset high value, e.g., as an expression of the scaled nominal profile $\beta$, here $\beta=2.35$. Conversely, for a less aggressive profile, we may set $\beta<1$ to match, e.g., $Q_{0.8}$ (of the 20% least deviating historical data).

### v2

# A practical scenario generation method for electricity prices on day-ahead and intraday spot markets

@@ Abstract
The increasing interest in demand-side management (DSM) as part of the energy cost optimization calls for effective methods to determine representative electricity prices for energy optimization and scheduling investigations. We propose a practical method to construct price profiles of day-ahead (DA) and intraday (ID) electricity spot markets. We construct single-day and single-week price profiles based on historical market time series to provide ready-to-use price data sets. Our method accounts for dominant mechanisms in price variation to preserve critical statistical features (e.g., mean and standard deviation) and transient patterns in the constructed profiles. Unlike common scenario generation approaches, the method is deterministic, with few degrees of freedom and minimal application effort. Our method ensures consistency between ID and DA price profiles when both are considered and introduces profile scaling to enable multiple scenario generation. Finally, we compare the constructed profiles to clustering techniques in a DSM case study, noting similar cost results.
@@
Published in Computers and Chemical Engineering: DOI:10.1016/j.compchemeng.2025.109118 © 2025. This manuscript version is made available under the CC-BY 4.0 license (https://creativecommons.org/licenses/by-nc-nd/4.0/).
@@ 1  Introduction
Optimization of energy costs has long been acknowledged as a topic of interest for the economic design and operation of chemical processes and energy systems (Friedler, 2010). Demand-side management (DSM) and demand response (DR) are energy-flexible operation paradigms that have have raised interest of industrial consumers to actively participate in the energy spot markets by adjustment of electricity demand (Pinson et al., 2014; Mitsos et al., 2018). With electricity prices on European DA spot markets having reached peak values as high as 4 000 €/MWh (Nord-Pool, 2023), the importance of industrial DSM is expected to grow further in the near future. Thus, economic studies on planning and scheduling of industrial processes necessitate a systematic way to select up-do-date electricity price scenarios. On the other hand, the process of selecting a price scenario should not distract from the main purpose of a scientific study on DSM. Thus, we herein propose a user-friendly intuitive approach to determine standard price profiles. Notice that we use the terms “price profiles” and “prices time series” interchangeably.
@@
Publications on DSM of industrial processes, including our own studies, such as Caspari et al. (2019); Mucci et al. (2023); El Wajeh et al. (2024), heavily rely on the chosen electricity price profiles that must represent current trends and allow for generalizable conclusions. Selecting a single historical profile or averaging historical data favors simplicity, and is essential in cases of computationally expensive calculations (e.g., Papadimitriou et al. (2023)), or a case study number that scales significantly with the number of price scenarios considered (e.g., Kelley et al. (2020)). Clearly, averaged price trends do not fully capture all features of historical profiles, as price averaging yields a single and exclusive price scenario. Additionally, data averaging does not account for the temporal evolution of prices to consider future price trends. Nonetheless, simple and sufficiently representative price profiles are usually adequate for evaluation of the theoretical potential of DR applications. Conversely, sophisticated scenario generation methods provide multiple price scenarios. These scenarios are rather valuable for active market participation, though less suitable for generalizable results on the considered DR method. Yet, these methods involve a more complex implementation, while forecasts of the long-term price trends are still subject to high uncertainty. Moreover, an extensive evaluation of all possible price scenarios, product demand situations, and other possible operating disturbances quickly results in a curse of dimensionality and is thus often beyond the scope of research studies on DR.
We propose a price scenario generation method, by considering two key price components of widely considered DR, namely the auction day-ahead (DA) and the real-time intraday (ID) electricity spot market. In order to decide on the time horizon of the generated price profiles, we note that in most related studies single-day or single-week price profiles are utilized, while less frequently longer horizons of historical data (e.g., up to one year) may be applied (Meese et al., 2016). The daily and weekly frequencies match the dominant patterns of the DA and ID price fluctuation identified in our previous studies (Schäfer et al. (2020); Germscheid et al. (2022)). We can thus extend the previously introduced averaging techniques to construct single-day and single-week DA and ID electricity price profiles.
@@
Herein, we aim to establish a simple method to construct meaningful and representative DA and ID ready-to-use price profiles to be used within future economic studies, e.g., cost optimal planning and scheduling, power-to-X, and energy-flexible design and operation. Within this method, we desire an option to allow consistency of the DA and ID prices, which is useful for DR across multiple electricity markets, e.g., in bidding strategies (Silva et al., 2022), or scheduling and control (Caspari et al., 2020). This ensures that the absolute DA and ID price values of a selected price scenario do not naïvely favor a certain DR strategy. An example of the effect of a systematic shift of DA and ID prices is given in our recent work Schulze et al. (2024), where flexible operation in the more volatile ID market appears less profitable than in the DA market, because of higher ID price values.
This property is not provided by existing methods, e.g., Maciejowska et al. (2019); Teichgraeber and Brandt (2019). The developed price profile will exhibit the following essential properties:
@@
    Consistency of DA and ID price profiles,
@@
To achieve the desired properties, we base our method on previous works on constructing average price profiles (Schäfer et al., 2020; Germscheid et al., 2022) and introduce adaptations tailored to the applications investigated. Because the focus of DR studies is on the economic potential of electricity price fluctuations rather than steady-state optimization, the price variance over the investigation period is assumed to hold greatest importance for optimal economic decision-making (Mathaba et al., 2014). Thus, our method allows for adjusting the magnitude of fluctuations in the price profile. Additionally, we ensure consistency of the DA and ID scenarios based on the near-zero mean distribution of the ID-DA deviation by equalizing the integrated DA and ID prices over the designated period. To underpin this approach, we provide statistical analyses of hourly DA and quarter-hourly real-time ID prices in historical data from European spot markets. Our approach can extend to other markets and price granularities (e.g., half-hourly) where similar observations hold.
@@
In the remainder of this article, we first present state-of-the-art approaches on electricity price scenario generation for DSM (Section 2). Subsequently, we introduce a method for constructing a practical 24-hour DA price profile from historical data of the last available 365 days (Section 3.1). Next, we utilize the same historical data to extract the ID price profile, based on ID - DA price deviations (Section 3.2). Then, we introduce a few modifications and apply the same approach to generate single-week (Section 4) practical profiles. We further create ready-to-use [^1] The open-access data sets are provided via Git (Link). DA and ID price profiles for future studies on DSM on the different time-scales presented and compare to historical data (Section 5). Last, we present a DR case study to contrast our method and state-of-the-art price generation techniques in terms of cost results (Section 6).
2  State of the art
Next, we review state-of-the-art techniques for electricity price scenario selection or generation. An overview of the presented methods is shown in Table 1, together with exemplary works of electricity price consideration. While not exhaustive, the presented classification covers the most common price generation methods and aligns with the objectives of our framework.
The most straightforward approach is to use recent historical profiles as scenarios. Several studies consider a single historical profile, e.g., 24 hours, that exhibits a typical behavior in terms of the price range and transient patterns (Dalle Ave et al., 2019). The authors then expect the findings to generalize well. Other works consider several different historical profiles commonly representing different extreme scenarios in their DSM application to highlight a range of the expected results (Caspari et al., 2020; Brée et al., 2019). Especially when studies on production design and planning are performed, a higher number of historical price profiles can be selected to capture the effect of different fluctuation patterns (Fürsch et al., 2013). The above scenario generation strategies are based on the use of raw historical data, yielding realistic results of the DSM application while keeping the number of scenarios low. However, they usually proceed ad-hoc, resulting in challenges for the representativity and generalizability of case study results.
A more sophisticated approach to determine characteristic periods in time-series data is time series aggregation (TSA) (Hoffmann et al., 2020). Within this category, the clustering of historical price data is a commonly used technique. Historical price clustering provides a set of price profiles that represent the entity of the historical data (Poncelet et al., 2017). All clustering techniques involve three core steps (Teichgraeber and Brandt, 2019). First, clustering criteria are defined to span the clustering space. For high-dimensional data vectors, e.g., price profiles, common low-dimensional clustering criteria include the mean and standard deviation of the data point (Wang et al., 2006). The vector of clustering criteria then represents the clustering candidates associated to original data vectors. Second, the distance between candidate cluster members and candidate cluster center is minimized to determine clusters and associated center points. Third, the cluster representative (characteristic profile) is derived from the cluster members.
The most common clustering strategies are $k$-means clustering, $k$-medoids clustering, and hierarchical clustering. A $k$-means clustering starts by randomly selecting $k$ cluster centers. Each cluster center is the mean, termed centroid, of the cluster members assigned to the cluster. Further, the cluster representative price profile is the mean of the profiles corresponding to the cluster (Jain, 2010). In $k$-medoids clustering, the most central cluster member, called medoid, is the center and derives the representative profile (Kaufman and Rousseeuw, 2009). Hierarchical clustering begins with single-member clusters that are iteratively merged using centroids as cluster centers. The cluster representatives are either defined as medoids or centroids (Nahmmacher et al., 2016; Teichgraeber and Brandt, 2019). Kotzur et al. (2018) compare different clustering techniques for price scenario generation applied to optimal process design. Furthermore, Wang et al. (2006) explore different clustering criteria to enhance clustering performance.
Although clustering reduces the number of scenarios compared to the original time series, multiple clusters within a DR study require multiple scheduling problems to be solved. Clustering techniques provide a more systematic way for historical profile selection compared to applying user-specified criteria with high degree of freedom. However, numerous options exist for configuring clustering methods. One such option is data scaling, which is essential for time-series clustering and can significantly impact results (Rager, 2015), but is not required for averaging and moment-matching techniques. Yet, the clustering results are only optimal with respect to the chosen criteria. In other words, the price profiles associated with cluster centers are the best representatives with respect to the chosen criteria, but they do not necessarily yield good results when applied to a DR case study. This can occur when employing too few criteria and risking loss of critical information. Conversely, using too many criteria can dilute specificity and produce profiles far from the average expected performance.
In an effort to enhance clustering performance, hybrid techniques that combine time-series feature extraction, such as principal component analysis (PCA), together with (machine learning) clustering techniques have been introduced for similar applications (e.g., load profiles) to capture underlying patterns and improve clustering (Almaimouni et al., 2018). Besides PCA, factor analysis methods are used to uncover latent factors and explain data variability. Examples include latent Dirichlet allocation (Chen et al., 2022) and autoencoder networks, e.g., variational autoencoders (Wang et al., 2022). Similarly, generative AI approaches like generative adversarial networks (Yilmaz et al., 2024) or jump-diffusion models (Meyer-Brandis and Tankov, 2008) can produce realistic data by learning the underlying data distribution. These methods can capture complex, nonlinear relationships and generate diverse, realistic profiles. However, they require increased computational complexity and careful tuning compared to moment-matching techniques.
An alternative way to produce practical price scenarios using historical data is to construct average price profiles over a historical period that respects fundamental characteristics of the spot market price profiles. Knittel and Roberts (2005) identify electricity price trends on different time scales highlighting recurring patterns and characteristic values of the price time series. Rahimiyan and Baringo (2016) perform a statistical analysis on historical DA and ID price data to quantify the price correlations. Röder et al. (2024) construct an average DA profile through permutation of the price elements of historical profile data together with averaging applicable to flexible operation for price peak avoidance. Schäfer et al. (2020) and Germscheid et al. (2022) utilize Fourier transforms to analyze principal frequencies of historical DA prices and ID-DA price deviations, respectively, identifying dominant modes at periods of 0.5 day and 1 day. By averaging one year of historical DA and ID price data, significant time-series frequencies, such as daily or weekly, are taken into account to construct a single profile. These periods correspond to significant price patterns and price fluctuations that can be exploited by most processes. The averaging method is straightforward to implement compared to clustering techniques and provides a single scenario. However, the approach may over-smooth the data and potentially obscure certain less prominent trends.
Finally, we mention probabilistic forecasting techniques, which exploit correlations of predicted parameters (e.g., weather forecasts) and are widely used to explain price behavior (Nowotarski and Weron, 2018). Following Monteiro et al. (2018) and Weron (2014), these methods can be categorized into prediction interval, density, and threshold forecasting. Univariate and multivariate probabilistic models are prevalent electricity price forecasting methods (Cuaresma et al., 2004; Rahimiyan and Baringo, 2016), but machine learning approaches like normalizing flows are also gaining attention (Cramer et al., 2023). Electricity prices forecasts enable DR computations in future applications, offering advantages over historical data, especially for long horizons and volatile markets (Nowotarski and Weron, 2018). Unlike data-averaging techniques, price forecasting methods are not restricted to representing data of the present and recent past. Instead, forecasting can incorporate trends of the future, although commonly restricted to the use of historical data and small forecasting windows (Alonso et al., 2016; Shah et al., 2021). Notably, however, both historical data based as well as complex scenario forecasting methods have been rather inaccurate at predicting the long-term evolution of electricity prices (Zareipour et al., 2010; Weron, 2014; Gabrielli et al., 2022), since this evolution is subject to policy changes and political events (Yang et al., 2017). At the same time, although practical for real-life application of DSM strategies, such as optimal bidding in a wholesale energy market (Zhang et al., 2020), price forecasts refer to a very specific time period in the future. They may therefore not provide a basis for generalizable conclusions when assessing the potential of an introduced DR case study.
In conclusion, we find all methods to be either fairly involved or subject to individual user decisions or randomness, i.e., non-deterministic. Hence, we aim to provide a simple yet systematic and deterministic method for constructing single-day and single-week electricity price scenarios based on historical time series data. Such a method shall involve few tuneable parameters and request little implementation effort. In the next section, we propose such a method based on averaging.
Table 1: Overview of the presented state-of-the-art techniques for electricity price scenario consideration. We distinguish between “historical” (real data) and “artificial” (synthetically generated) profiles, “multiple” or “single” number of generated profiles, and “past” or “future” represented time windows or price realization.
Technique Key Derived Derived
features price profile References
datasets window
Recent prices price data historical, past Dalle Ave et al. (2019),
qualitative/ single/ Caspari et al. (2020),
quantitative multiple Brée et al. (2019),
selection Fürsch et al. (2013)
Clustering $k$-means, historical/ past Kotzur et al. (2018),
$k$-medoids, artificial, Poncelet et al. (2017),
hierarchical multiple Teichgraeber and Brandt (2019)
Hybrid feature historical/ past Almaimouni et al. (2018)
clustering extraction artificial, Yilmaz et al. (2024)
& clustering multiple Meyer-Brandis and Tankov (2008)
Averaging of one-year artificial, past Röder et al. (2024)
historical single Germscheid et al. (2022)
data Schäfer et al. (2020)
prediction future Nowotarski and Weron (2018)
interval, Monteiro et al. (2018)
Probabilistic density, artificial, Weron (2014); Zhang et al. (2020)
forecasting threshold, single Cramer et al. (2023)
univariate, Cuaresma et al. (2004)
multivariate Rahimiyan and Baringo (2016)
@@ 3  Single-day price profile
In this section, we present the methodology for constructing representative single-day DA (Section 3.1) and ID (Section 3.2) price profiles using a full-year time-series data set.
@@ 3.1 DA price profile
We begin with the construction of the DA profile of one representative day. To this end, we select a price data set of 365 consecutive days to exclude seasonal effects (Knittel and Roberts, 2005). The 365 days horizon may be chosen arbitrarily and if desired extended to multiple years. We remark that incidents and trends contained in the price data set, e.g., caused by political events and developments, will affect the price scenario created. Here, we use up-to-date price values of the full year 2023. Similar to our prior works (Schäfer et al., 2020; Germscheid et al., 2022), we first average the historical DA prices for every day hour, k=1,2,,24k=1,2,...,24, individually [^2] Notice that for $k\geq 2$, Eq. (1) may be equivalently rewritted as $\overline{DA}_{k}=\overline{DA}_{k-1}+\overline{\Delta DA}_{k}$, where $\overline{\Delta DA}_{k}=1/N\sum_{i=1}^{N}{(DA_{i,k}-DA_{i,k-1})}$. Hence, this profile captures the moving average price development over a day. :
@@
where $N=365$ is the number of days of selected historical data, $DA_{i,k}$ is the DA price at hour $k$ of day i=1,2,,Ni=1,2,...,N, and $\overline{DA}_{k}$ is the average DA price of hour $k$. If desired, days with significant outliers, weekends or holidays may be excluded. However, here we consider the full year.
@@
Fig. 1: Average (Unscaled) and scaled average ($\beta=1.47$) (Nominal) 24-hour DA price profile generated from historical 2023 data (EPEX, 2023).
@@
Clearly, setting $\beta=1$ corresponds to no scaling and practically represents the centroid profile derived by any clustering technique which applies a single cluster and derivation of the representative profiles with centroids. The resulting scenario exhibits low deviating behavior according to the historical data (Fig. 2). It could thus serve as a price scenario of moderate fluctuations. Such an averaging-out of the data variation is also observed in centroid clustering methods, like $k$-means clustering (Kotzur et al., 2018; Schütz et al., 2018). To minimize scenario complexity, we can thus use the $\beta=1$-profile as an Unscaled scenario. Importantly, the mean value $\overline{DA}$ remains identical irrespective of scaling.
Therefore, as the daily standard deviation of the average profile does not represent the daily standard deviation of all data due to smoothing of the data (Schäfer et al., 2020), we use $\beta$ to correct the averaged results. This allows to preserve a key time-series property critical for DR applications, namely the magnitude of price deviation. Thus, in order to obtain a Nominal ready-to-use profile, we choose $\beta$ such that:
@@
For example, for the full year 2023 and EPEX spot market, we receive the value $\beta=1.47$. The resulting DA profile is illustrated by Fig. 1. Here, the characteristic double peak profile is visible with minimal values occurring during the 4th and 14th hour as well as maximal values during the 8th and the 19th. The absolute price range after scaling ($\beta=1.47$) is $103.46\,\mathrm{€/MWh}$ with a mean value of $\overline{DA}=95.18\,\mathrm{€/MWh}$.
If a more distinct price variation is desired, we can specify a larger value of $\beta$. In order to systematically choose a larger $\beta$ factor that corresponds to some extreme behavior of the historical data (Requirement 2), we represent the density distribution of the standard deviations $\sigma(DA_{i,1},...,DA_{i,24})$ across all days i=1,2,,Ni=1,2,...,N through a histogram, Fig. 2. Subsequently, we identify a quantile indicative of high fluctuations within the historical data. We exemplary show the quantile 3.4 ($Q_{3.4}$), corresponding to the 15% highest values of the daily standard deviations, Fig. 2. The value of the quantile is the corresponding standard deviation, namely:
@@
giving $\beta=1.97$ for the full year 2023. Other options could include the choice of a preset high value, e.g., as an expression of the scaling parameter $\beta$ that corresponds to the Nominal profile. Conversely, for a less aggressive profile, we may set $\beta<1$ to match, e.g., $Q_{0.8}$ (of the 20% least deviating historical data). The use of such variations on the averaged profile variance enables various different scenarios to be used in a DR case study. Examples are optimal design and scheduling, where a sensitivity analysis over the prices allows to determine limits in equipment oversizing (e.g., Röder et al. (2024)).